OpenStudy (anonymous):

Find variation constant in which y varies as x and following conditions y=6 when x=1/3 k=?

myininaya (myininaya):

varies directly or inversely?

OpenStudy (anonymous):

inversley

myininaya (myininaya):

y=k/x

myininaya (myininaya):

6=k/(1/3) 6=3k k=2

OpenStudy (anonymous):

ah .. now I see thanks

myininaya (myininaya):

:) if we had directly it would be y=kx inversely think division directly think multiplication

OpenStudy (anonymous):

oh ok.. ya I had some directly ones, now I can think of those in multiplication..that helps

myininaya (myininaya):

but dont forget about your constant

OpenStudy (anonymous):

ok..:)